OpenStudy (anonymous):

Solve for x: x/4+6=10

OpenStudy (anonymous):

use transposition :)

OpenStudy (anonymous):

x/4 + 6 = 10 x/4 = 4 do the next step x=?

OpenStudy (anonymous):

Wouldnt u square root it?

OpenStudy (anonymous):

x/4=10-6 x=4(10-6) x=? now you find x :)

OpenStudy (anonymous):

no you wouldn't

OpenStudy (anonymous):

no! you wont haha coz it has no square or x^2 :D

OpenStudy (anonymous):

Oh lol

OpenStudy (austinl):

\[\frac{x}{4}+6=10\] Is this correct? If so we would subtract the 6 over first to get \[\frac{x}{4}=4\] Then multiply by 4 to arrive at \[x=16\]

OpenStudy (anonymous):

Ahhh 16 x=16

OpenStudy (austinl):

x=16 yes